# Qwen2.5-Coder Technical Report

### v1

@@ Abstract
In this report, we introduce the Qwen2.5-Coder series, a significant upgrade from its predecessor, CodeQwen1.5. This series includes two models: Qwen2.5-Coder-1.5B and Qwen2.5-Coder-7B. As a code-specific model, Qwen2.5-Coder is built upon the Qwen2.5 architecture and continues pretrained on a vast corpus of over 5.5 trillion tokens. Through meticulous data cleaning, scalable synthetic data generation, and balanced data mixing, Qwen2.5-Coder demonstrates impressive code generation capabilities while retaining general versatility. The model has been evaluated on a wide range of code-related tasks, achieving state-of-the-art (SOTA) performance across more than 10 benchmarks, including code generation, completion, reasoning, and repair, consistently outperforming larger models of the same model size. We believe that the release of the Qwen2.5-Coder series will not only push the boundaries of research in code intelligence but also, through its permissive licensing, encourage broader adoption by developers in real-world applications.
@@ 1 Introduction
With the rapid development of large language models (LLMs) (Brown, 2020; Achiam et al., 2023; Touvron et al., 2023; Dubey et al., 2024; Jiang et al., 2023; Bai et al., 2023b; Yang et al., 2024; Anthropic, 2024; OpenAI, 2024), code-specific language models have garnered significant attention in the community. Built upon pretrained LLMs, code LLMs such as the StarCoder series (Li et al., 2023; Lozhkov et al., 2024), CodeLlama series (Roziere et al., 2023), DeepSeek-Coder series (Guo et al., 2024), CodeQwen1.5 (Bai et al., 2023a), and CodeStral (team, 2024), have demonstrated superior performance in coding evaluations (Chen et al., 2021; Austin et al., 2021; Cassano et al., 2022; Jain et al., 2024). However, in comparison with the recently state-of-the-art proprietary LLMs, Claude-3.5-Sonnet (Anthropic, 2024) and GPT-4o (OpenAI, 2024), the code LLMs are still falling behind, either open-source or proprietary models.
Building upon our previous work, CodeQwen1.5, we are excited to introduce Qwen2.5-Coder, a new series of language models designed to achieve top-tier performance in coding tasks at various model sizes. Qwen2.5-Coder models are derived from the Qwen2.5 LLMs, inheriting their advanced architecture and tokenizer. These models are pretrained on extensive datasets and further fine-tuned on carefully curated instruction datasets specifically designed for coding tasks. The series includes models with 1.5B and 7B parameters, along with their instruction-tuned variants. We are committed to fostering research and innovation in the field of code LLMs, coding agents, and coding assistant applications. Therefore, we are open-sourcing the Qwen2.5-Coder models to the community to support and accelerate advancements in these areas.
@@
This report introduces the Qwen2.5-Coder series, an upgraded version of CodeQwen1.5, featuring two models: Qwen2.5-Coder-1.5B and Qwen2.5-Coder-7B. Built on the Qwen2.5 architecture and pretrained on over 5.5 trillion tokens, these code-specific models demonstrate exceptional code generation capabilities while maintaining general versatility. Through rigorous data processing and training techniques, Qwen2.5-Coder achieves state-of-the-art performance across more than 10 code-related benchmarks, outperforming larger models in various tasks. The release of these models aims to advance code intelligence research and promote widespread adoption in real-world applications, facilitated by permissive licensing.
@@ 2 Model Architecture
The architecture of Qwen2.5-Coder is the same as Qwen2.5. Table 1 shows the architecture of Qwen2.5-Coder for two different model sizes: 1.5B and 7B parameters. Both sizes share the same architecture in terms of layers, having 28 layers and a head size of 128. However, they differ in several key aspects. The 1.5B model has a hidden size of 1,536, while the 7B model has a much larger hidden size of 3,584. The 1.5B model uses 12 query heads and 2 key-value heads, whereas the 7B model uses 28 query heads and 4 key-value heads, reflecting its larger capacity. The intermediate size also scales with model size, being 8,960 for the 1.5B model and 18,944 for the 7B model. Additionally, the 1.5B model employs embedding tying, while the 7B model does not. Both models share the same vocabulary size of 151,646 tokens and have been trained on 5.5 trillion tokens.
@@
Configuration Qwen2.5-Coder 1.5B Qwen2.5-Coder 7B
Hidden Size 1,536 3,584
# Layers 28 28
# Query Heads 12 28
# KV Heads 2 4
Head Size 128 128
Intermediate Size 8,960 18,944
Embedding Tying True False
Vocabulary Size 151,646 151,646
# Trained Tokens 5.5T 5.5T
@@ 3.1 Pretraining Data
Large-scale, high-quality, and diverse data forms the foundation of pre-trained models. To this end, we constructed a dataset named Qwen2.5-Coder-Data. This dataset comprises five key data types: Source Code Data, Text-Code Grounding Data, Synthetic Data, Math Data, and Text Data. In this section, we provide a brief overview of the sources and cleaning methods applied to these datasets.
@@ 3.1.1 Data Composition
We collected public repositories from GitHub created before February 2024, spanning 92 programming languages. Similar to StarCoder2 (Lozhkov et al., 2024) and DS-Coder (Guo et al., 2024), we applied a series of rule-based filtering methods. In addition to raw code, we also collected data from Pull Requests, Commits, Jupyter Notebooks, and Kaggle datasets, all of which were subjected to similar rule-based cleaning techniques.
@@
In Qwen2.5-Coder, we applied this process iteratively. As shown in Figure 1, each iteration resulted in improvement. Through 4-stage filtering, the average scores on HumanEval and MBPP increased from 41.6% to 46.8% compared to the baseline, demonstrating the value of high-quality Text-Code Grounding Data for code generation.
@@ 3.1.2 Data Mixture
Balancing Code, Math, and Text data is crucial for building a robust foundational model. Although the research community has explored this balance before, there is limited evidence regarding its scalability to large datasets. To address this, we conducted empirical experiments with different ratios of Code, Math, and Text data, designing multiple experiments to identify an optimal combination rapidly. Specifically, as shown in Table 3, we compared three different Code: Text ratios — 100:0:0, 85:10:5, and 70:20:10.
@@ 3.2.2 Repo-Level Pretraining
After file-level pretraining, we turn to repo-level pretraining, aimed at enhancing the model’s long-context capabilities. In this stage, the context length is extended from 8,192 tokens to 32,768 tokens, and RoPE’s base frequency is adjusted from 10,000 to 1,000,000. To further leverage the model’s extrapolation potential, we applied the YARN mechanism (Peng et al., 2023), enabling the model to handle sequences up to 131,072 (132K) tokens.
In this stage, we used a large amount of high-quality, long-code data ($\approx$ 300B) and extended file-level FIM to the repo-level FIM followed by methods described in Lozhkov et al. (2024), with the specific format shown in Figure 4.
@@ 4.1 A Recipe for Instruction Data
We fine-tune a CodeBERT to perform the language identification model to categorize documents into nearly 100 programming languages. We keep the instruction data of the mainstream programming languages and randomly discard a portion of the instruction data of the long-tail languages. If a given sample contains very little code data or even no code snippets, the sample will possibly be classified into “No Programming Language” tag. We remove most of the samples without code snippets to keep the code generation capability of our instruction model.
@@
For the unsupervised data (code snippets) massively existing in many websites (e.g. GitHub), we try to construct the supervised instruction dataset. Specifically, we use the LLM to generate the instruction from the code snippets within 1024 tokens and then we use the code LLM to generate the response Sun et al. (2024). Finally, we use the LLM scorer to filter the low-quality ones to obtain the final pair. Given the code snippets of different programming languages, we construct an instruction dataset from the code snippets. To increase the diversity of the instruction dataset. Conversely, we first generate the answers from the code. Then we use the LLM scorer to filter the low-quality to obtain the final triplet. Similarly, given the code snippets of different programming languages, we can construct an instruction dataset with the universal code from the code snippets. To fully unleash the potential of our proposed method, we also include the open-source instruction dataset in the seed instruction dataset. Finally, we combine three parts instruction dataset for supervised fine-tuning.
@@
To completely evaluate the quality of the created instruction pair, we introduce several scoring points for each sample: (1) Question&Answer Consistency: Whether Q&A are consistent and correct for fine-tuning. (2) Question&Answer Relevance: Whether Q&A are related to the computer field. (3) Question&Answer Difficulty: Whether Q&A are sufficiently challenging. (4) Code Exist: Whether the code is provied in question or answer. (5) Code Correctness: Evaluate whether the provided code is free from syntax errors and logical flaws. (6) Consider factors like proper variable naming, code indentation, and adherence to best practices. (7) Code Clarity: Assess how clear and understandable the code is. Evaluate if it uses meaningful variable names, proper comments, and follows a consistent coding style. (8) Code Comments: Evaluate the presence of comments and their usefulness in explaining the code’s functionality. (9) Easy to Learn: determine its educational value for a student whose goal is to learn basic coding concepts. After gaining all scores $(s_{1},\dots,s_{n})$, we can get the final score with $s=w_{1}s_{1}+\dots+w_{n}s_{n}$, where $(w_{1},\dots,w_{n})$ are a series of pre-defined weights.
@@ 5 Decontamination
To ensure that Qwen2.5-Coder does not produce inflated results due to test set leakage, we performed decontamination on all data, including both pre-training and post-training datasets. We removed key datasets such as HumanEval, MBPP, GSM8K, and MATH. The filtering was done using a 10-gram overlap method, where any training data with a 10-gram string-level overlap with the test data was removed.
@@ 6 Evaluation on Base Models
Qwen2.5-Coder-1.5B https://hf.co/qwen/Qwen/Qwen2.5-Coder-1.5B
Qwen2.5-Coder-7B https://hf.co/qwen/Qwen/Qwen2.5-Coder-7B
CodeQwen1.5-7B-Base https://huggingface.co/Qwen/CodeQwen1.5-7B
@@ 6.1 Code Generation
DS-Coder-V2-Lite-Base 2.4/16B 40.9 34.1 71.9 59.4 - 30.6 8.1
CodeQwen1.5-7B-Base 7B 51.8 45.7 72.2 60.2 61.8 45.6 15.6
Qwen2.5-Coder-7B-Base 7B 61.6 53.0 76.9 62.9 68.8 45.8 16.2
@@
CodeQwen1.5-7B-Base 7B 51.8 52.2 42.4 46.6 52.2 55.7 36.7 49.7 48.4
Qwen2.5-Coder-7B-Base 7B 61.6 62.1 53.2 59.0 64.2 60.8 38.6 60.3 57.5
@@
As shown in Table 5, Qwen2.5-Coder have shown impressive performance in basic code generation, achieving state-of-the-art results among open-source models of the same size and surpassing even larger models. In particular, Qwen2.5-Coder-7B-Base outperforms the previous best dense model, DS-Coder-33B-Base, across all five metrics.
@@
The evaluations mentioned above focus on the Python language. However, we expect a strong code model to be not only proficient in Python but also versatile across multiple programming languages to meet the complex and evolving demands of software development. To more comprehensively evaluate Qwen2.5-Coder’s proficiency in handling multiple programming languages, we selected the MultiPL-E (Cassano et al., 2022) and choose to evaluate eight mainstream languages from these benchmark, including Python, C++, Java, PHP, TypeScript, C#, Bash and JavaScript.
@@ 6.2 Code Completion
HumanEval Infilling
Many developer aid tools rely on the capability to autocomplete code based on preceding and succeeding code snippets. Qwen2.5-Coder utilizes the Fill-In-the-Middle (FIM) training strategy, as introduced in Bavarian et al. (2022), enabling the model to generate code that is contextually coherent. To assess its code completion proficiency, we utilize the HumanEval Infilling benchmark Allal et al. (2023). This benchmark challenges the model to accurately predict missing sections of code within tasks derived from HumanEval. We use the single-line infilling settings across Python, Java, and JavaScript, focusing on predicting a single line of code within given contexts. Performance was measured using the Exact Match metric, which determines the proportion of the first generated code line that precisely match the ground truth.
The table 7 illustrates that Qwen2.5-Coder surpasses alternative models concerning model size. Specifically, Qwen2.5-Coder-1.5B achieves an average performance improvement of 3.7%, rivaling the majority of models exceeding 6 billion parameters. Moreover, Qwen2.5-Coder-7B-Base stands as the leading model among those over 6 billion parameters, matching the performance of the formidable 33 billion parameter model, DS-Coder-33B-Base. Notably, we excluded DS-Coder-v2-236B from comparison due to its design focus not being on code completion tasks.
Model Size HumanEval-Infilling
@@
StarCoder2-3B 3B 70.9 84.4 81.8 79.0
DS-Coder-1.3B-Base 1.3B 72.8 84.3 81.7 79.6
Qwen2.5-Coder-1.5B 1.5B 77.0 85.6 85.0 82.5
@@
StarCoder2-7B 7B 70.8 86.0 84.4 80.4
StarCoder2-15B 15B 78.1 87.4 84.1 81.3
DS-Coder-6.7B-Base 6.7B 78.1 87.4 84.1 83.2
DS-Coder-V2-Lite-Base 2.4/16B 78.7 87.8 85.9 84.1
CodeQwen1.5-7B-Base 7B 75.8 85.7 85.0 82.2
Qwen2.5-Coder-7B-Base 7B 79.7 88.5 87.6 85.3

### v2

@@ Abstract
In this report, we introduce the Qwen2.5-Coder series, a significant upgrade from its predecessor, CodeQwen1.5. This series includes six models: Qwen2.5-Coder-(0.5B/1.5B/3B/7B/14B/32B). As a code-specific model, Qwen2.5-Coder is built upon the Qwen2.5 architecture and continues pretrained on a vast corpus of over 5.5 trillion tokens. Through meticulous data cleaning, scalable synthetic data generation, and balanced data mixing, Qwen2.5-Coder demonstrates impressive code generation capabilities while retaining general and math skills. These models have been evaluated on a wide range of code-related tasks, achieving state-of-the-art (SOTA) performance across more than 10 benchmarks, including code generation, completion, reasoning, and repair, consistently outperforming larger models of the same model size. We believe that the release of the Qwen2.5-Coder series will advance research in code intelligence and, with its permissive licensing, support wider adoption by developers in real-world applications.
@@ 1 Introduction
With the rapid development of large language models (LLMs) (Brown, 2020; Achiam et al., 2023; Touvron et al., 2023; Dubey et al., 2024; Jiang et al., 2023; Bai et al., 2023; Yang et al., 2024; Anthropic, 2024; OpenAI, 2024), code-specific language models have garnered significant attention in the community. Built upon pre-trained LLMs, code LLMs such as the StarCoder series (Li et al., 2023; Lozhkov et al., 2024), CodeLlama series (Roziere et al., 2023), DeepSeek-Coder series (Guo et al., 2024a), CodeQwen1.5 (Qwen, 2024), and CodeStral (MistralAI, 2024), have demonstrated superior performance in coding evaluations (Chen et al., 2021; Austin et al., 2021; Cassano et al., 2022; Jain et al., 2024; Liu et al., 2024a; Li et al., 2024b; Guo et al., 2024b; Wu et al., 2024b). However, in comparison with the recently state-of-the-art proprietary LLMs, Claude-3.5-Sonnet (Anthropic, 2024) and GPT-4o (OpenAI, 2024), the code LLMs are still falling behind, either open-source or proprietary models.
Building upon our previous work, CodeQwen1.5, we are excited to introduce Qwen2.5-Coder, a new series of language models designed to achieve top-tier performance in coding tasks at various model sizes. Qwen2.5-Coder models are derived from the Qwen2.5 LLMs, inheriting their advanced architecture and tokenizer. These models are trained on extensive datasets and further fine-tuned on carefully curated instruction datasets specifically designed for coding tasks. We are committed to fostering research and innovation in the field of code LLMs, coding agents, and coding assistant applications. Therefore, we release the Powerful, Diverse, and Practical Qwen2.5-Coder series, dedicated to continuously promoting the development of Open CodeLLMs. (1) Powerful: Qwen2.5-Coder-32B-Instruct has become the current SOTA open-source code model, matching the coding capabilities of GPT-4o. While demonstrating strong and comprehensive coding abilities, it also possesses good general and mathematical skills. (2) Diverse: Qwen2.5-Coder series brings six model sizes, including 0.5B/1.5B/3B/7B/14B/32B. Qwen2.5-Coder has covered six mainstream model sizes to meet the needs of different developers. (3) Practical: We explore the practicality of Qwen2.5-Coder in two scenarios, including code assistants and Artifacts, with some examples showcasing the potential applications of Qwen2.5-Coder in real-world scenarios
@@
To evaluate the effectiveness of Qwen2.5-Coder, we conducted an extensive evaluation on a suite of popular benchmarks. The results highlight Qwen2.5-Coder’s superior code generation capabilities, achieving state-of-the-art performance across more than ten code-focused benchmarks while maintaining robust general and mathematical reasoning abilities. This model outperforms larger code models on a variety of tasks. The release of these models aims to advance code intelligence research and promote widespread adoption in real-world applications, facilitated by permissive licensing.
@@ 2 Model Architecture
The architecture of Qwen2.5-Coder is derived directly from Qwen2.5. Table 1 outlines the architecture of Qwen2.5-Coder across six different model sizes: 0.5B, 1.5B, 3B, 7B, 14B, and 32B parameters. While all sizes share the same architecture in terms of head size, they differ in several other key aspects. With exceptions like the 1.5B model having a larger intermediate size and the 3B model having more layers, most parameters generally increase as the model size scales up. Comparing the 7B and 32B models for instance: the 7B model features a hidden size of 3,584, whereas the 32B model has a hidden size of 5,120. The 7B model uses 28 query heads and 4 key-value heads, while the 32B model uses 40 query heads and 8 key-value heads, reflecting its enhanced capacity. Similarly, the intermediate size scales with model size, being 18,944 for the 7B model and 27,648 for the 32B model. Additionally, smaller models use embedding tying, while larger models do not. Both models have a vocabulary size of 151,646 tokens and are trained on 5.5 trillion tokens.
@@
Configuration 0.5B 1.5B 3B 7B 14B 32B
Hidden Size 896 1,536 2048 3,584 5120 5120
# Layers 24 28 36 28 48 64
# Query Heads 14 12 16 28 40 40
# KV Heads 2 2 2 4 8 8
Head Size 128 128 128 128 128 128
Intermediate Size 4,864 8,960 4,864 18,944 13824 27648
Embedding Tying
Vocabulary Size 151,646 151,646 151,646 151,646 151,646 151,646
# Trained Tokens 5.5T 5.5T 5.5T 5.5T 5.5T 5.5T
@@ 3.1 Pretraining Data
Large-scale, high-quality, and diverse data forms the foundation of pre-trained models. To this end, we constructed a dataset named Qwen2.5-Coder-Data. This dataset comprises five key data types: Source Code Data, Text-Code Grounding Data, Synthetic Data, Math Data and Text Data. In this section, we provide a brief overview of the sources and cleaning methods applied to these datasets.
@@ 3.1.1 Data Composition
We collected public repositories from GitHub created before February 2024, spanning 92 programming languages. Similar to StarCoder2 (Lozhkov et al., 2024) and DS-Coder (Guo et al., 2024a), we applied a series of rule-based filtering methods. In addition to raw code, we also collected data from Pull Requests, Commits, Jupyter Notebooks, and Kaggle datasets, all of which were subjected to similar rule-based cleaning techniques.
@@
In Qwen2.5-Coder, we applied this process iteratively. As shown in Figure 1, each iteration resulted in improvement for Qwen2.5-Coder-1.5B. Through 4-stage filtering, the average scores on HumanEval and MBPP increased from 41.6% to 46.8% compared to the baseline, demonstrating the value of high-quality Text-Code Grounding Data for code generation.
@@ 3.1.2 Data Mixture
Balancing Code, Math, and Text data is crucial for building a foundational model. Although the research community has explored this balance before, there is limited evidence regarding its scalability to large datasets. To address this, we conducted empirical experiments with different ratios of Code, Math, and Text data, designing multiple experiments to identify an optimal combination rapidly. Specifically, as shown in Table 3, we compared three different Code for Qwen2.5-Coder-7B: Text ratios — 100:0:0, 85:10:5, and 70:20:10.
@@ 3.2.2 Repo-Level Pretraining
After file-level pretraining, we turn to repo-level pretraining, aimed at enhancing the model’s long-context capabilities. In this stage, the context length is extended from 8,192 tokens to 32,768 tokens, and RoPE’s base frequency is adjusted from 10,000 to 1,000,000. To further leverage the model’s extrapolation potential, we applied the YARN mechanism (Peng et al., 2023), enabling the model to handle sequences up to 131,072 (128K) tokens.
In this stage, we used a large amount of high-quality, long-context code data ($\approx$ 300B) and extended file-level FIM to the repo-level FIM followed by methods described in Lozhkov et al. (2024), with the specific format shown in Figure 4.
@@ 4.1 A Recipe for Instruction Data
We fine-tune a CodeBERT (Feng et al., 2020) to perform the language identification model to categorize documents into nearly 100 programming languages. We keep the instruction data of the mainstream programming languages and randomly discard a portion of the instruction data of the long-tail languages. If a given sample contains very little code data or even no code snippets, the sample will possibly be classified into “No Programming Language” tag. Since too many instruction samples without code snippets hurt the model performance on code generation tasks (e.g. MultiPL-E, McEval, and MdEval), we remove most of the samples without code snippets to keep the code generation capability of our instruction model.
@@
For the unsupervised data (code snippets) massively existing in many websites (e.g. GitHub), we try to construct the supervised instruction dataset using LLM. Specifically, we use the LLM to generate the instruction from the code snippets within 1024 tokens and then we use the code LLM to generate the response (Wei et al., 2024; Sun et al., 2024; Yu et al., 2024). Finally, we use the LLM scorer to filter the low-quality ones to obtain the final pair. Given the code snippets of different programming languages, we construct an instruction dataset from the code snippets. To fully unleash the potential of our proposed method, we also include the open-source instruction dataset (e.g. McEval-Instruct for massively multilingual code generation and debugging [^1] https://huggingface.co/datasets/Multilingual-Multimodal-NLP/McEval-Instruct) in the seed instruction dataset. Finally, we combine the instruction data from the GitHub code snippet and open-source instructions for supervised fine-tuning.
@@
To completely evaluate the quality of the created instruction pair, we introduce several scoring points for each sample: (1) Question&Answer Consistency: Whether Q&A are consistent and correct for fine-tuning. (2) Question&Answer Relevance: Whether Q&A are related to the computer field. (3) Question&Answer Difficulty: Whether Q&A are sufficiently challenging. (4) Code Exist: Whether the code is provided in question or answer. (5) Code Correctness: Evaluate whether the provided code is free from syntax errors and logical flaws. (6) Consider factors like proper variable naming, code indentation, and adherence to best practices. (7) Code Clarity: Assess how clear and understandable the code is. Evaluate if it uses meaningful variable names, proper comments, and follows a consistent coding style. (8) Code Comments: Evaluate the presence of comments and their usefulness in explaining the code’s functionality. (9) Easy to Learn: determine its educational value for a student whose goal is to learn basic coding concepts. After gaining all scores $(s_{1},\dots,s_{n})$, we can get the final score with $s=w_{1}s_{1}+\dots+w_{n}s_{n}$, where $(w_{1},\dots,w_{n})$ are a series of pre-defined weights.
@@ 4.2 Training Policy
Direct Preference Optimization for Code
After obtaining the SFT model, we further align the Qwen2.5-Coder with the help of offline direct preference optimization (DPO) (Rafailov et al., 2023). Given that human feedback is highly labor-intensive, we use a multilingual code sandbox to provide code execution feedback, while an LLM is utilized for human judgment feedback. For the algorithm-like and self-contained code snippets, we generate the test cases to check the correctness of the code as the code execution feedback, including Python, Java, and other languages. For other complex code snippets, we use LLM-as-a-judge (Zheng et al., 2023) to decide which code snippet is better. Further, we combine the code DPO data and common data for offline DPO training.
@@ 5 Decontamination
To ensure that Qwen2.5-Coder does not produce inflated results due to test set leakage, we performed decontamination on all data, including both pre-training and post-training datasets. We removed key datasets such as HumanEval, MBPP, GSM8K, and MATH. The filtering was done using a 10-gram overlap method, where any training data with a 10-gram word-level overlap with the test data was removed.
@@ 6 Evaluation on Base Models
Qwen2.5-Coder-0.5B https://hf.co/Qwen/Qwen2.5-Coder-0.5B
Qwen2.5-Coder-1.5B https://hf.co/Qwen/Qwen2.5-Coder-1.5B
Qwen2.5-Coder-3B https://hf.co/Qwen/Qwen2.5-Coder-3B
Qwen2.5-Coder-7B https://hf.co/Qwen/Qwen2.5-Coder-7B
Qwen2.5-Coder-14B https://hf.co/Qwen/Qwen2.5-Coder-14B
Qwen2.5-Coder-32B https://hf.co/Qwen/Qwen2.5-Coder-32B
CodeQwen1.5-7B https://hf.co/Qwen/CodeQwen1.5-7B
@@ 6.1 Code Generation
DS-Coder-V2-Lite-Base 2.4/16B 40.9 34.1 71.9 59.4 62.6 30.6 8.1
CodeQwen1.5-7B 7B 51.8 45.7 72.2 60.2 61.8 45.6 15.6
Qwen2.5-Coder-7B 7B 61.6 53.0 76.9 62.9 68.8 45.8 16.2
14B+ Models
StarCoder2-15B 15B 46.3 37.8 66.2 53.1 57.0 38.4 12.2
Qwen2.5-Coder-14B 14B 64.0 57.9 81.0 66.7 71.4 51.8 22.3
@@
CodeQwen1.5-7B 7B 51.8 52.2 42.4 46.6 52.2 55.7 36.7 49.7 48.4
Qwen2.5-Coder-7B 7B 61.6 62.1 53.2 59.0 64.2 60.8 38.6 60.3 57.5
14B+ Models
StarCoder2-15B 15B 46.3 47.2 46.2 39.1 42.1 53.2 15.8 43.5 41.7
Qwen2.5-Coder-14B 14B 64.0 69.6 46.8 64.6 69.2 63.3 39.9 61.5 59.9
@@
As shown in Table 5, Qwen2.5-Coder have shown impressive performance in basic code generation, achieving state-of-the-art results among open-source models of the same size and surpassing even larger models. In particular, Qwen2.5-Coder-7B outperforms the previous best dense model, DS-Coder-33B, across all five metrics.
@@
The evaluations mentioned above focus on the Python language. However, we expect a strong code model to be not only proficient in Python but also versatile across multiple programming languages to meet the complex and evolving demands of software development. To more comprehensively evaluate Qwen2.5-Coder’s proficiency in handling multiple programming languages, we selected the MultiPL-E (Cassano et al., 2022) and chose to evaluate eight mainstream languages from this benchmark, including Python, C++, Java, PHP, TypeScript, C#, Bash and JavaScript.
@@ 6.2 Code Completion
Many developer aid tools rely on the capability to autocomplete code based on preceding and succeeding code snippets. Qwen2.5-Coder utilizes the Fill-In-the-Middle (FIM) training strategy, as introduced in Bavarian et al. (2022), enabling the model to generate code that is contextually coherent. To assess its code completion proficiency, we utilize the HumanEval-FIM benchmark (Allal et al., 2023), CrossCodeEval (Ding et al., 2024), CrossCodeLongEval (Wu et al., 2024a), RepoEval (Zhang et al., 2023) and SAFIM (Gong et al., 2024). Figure 5 shows the overall evaluation results of Qwen2.5-Coder-32B on different code completion benchmarks.
Figure 5: The code completion performance of competitive models on five benchmarks, Humaneval-FIM, SAFIM, CrossCodeEval, RepoEval, CrossCodeLongEval.
Humaneval-FIM benchmark challenges the model to accurately predict missing sections of code within tasks derived from Humaneval. We use the single-line infilling settings across Python, Java, and JavaScript, focusing on predicting a single line of code within given contexts. Performance was measured using the Exact Match metric, which determines the proportion of the first generated code line that precisely match the ground truth. The table 7 illustrates that Qwen2.5-Coder surpasses alternative models concerning model size. Specifically, Qwen2.5-Coder-1.5B achieves an average performance improvement of 3.7%, rivaling the majority of models exceeding 6 billion parameters. Moreover, Qwen2.5-Coder-7B stands as the leading model among those over 6 billion parameters, matching the performance of the formidable 33 billion parameter model, DS-Coder-33B-Base. Notably, we excluded DS-Coder-v2-236B from comparison due to its design focus not being on code completion tasks.
Model Size Humaneval-FIM
@@
DS-Coder-1.3B-Base 1.3B 72.8 84.3 81.7 80.7
Qwen2.5-Coder-1.5B 1.5B 77.0 85.6 85.0 83.5
3B+ Models
StarCoder2-3B 3B 70.9 84.4 81.8 80.4
Qwen2.5-Coder-3B 3B 78.7 88.0 87.4 85.7
@@
StarCoder2-7B 7B 70.8 86.0 84.4 82.0
DS-Coder-6.7B-Base 6.7B 78.1 87.4 84.1 84.0
DS-Coder-V2-Lite-Base 2.4/16B 78.7 87.8 85.9 85.0
CodeQwen1.5-7B 7B 75.8 85.7 85.0 83.3
Qwen2.5-Coder-7B 7B 79.7 88.5 87.6 86.2
14B+ Models
StarCoder2-15B 15B 74.2 85.2 84.6 82.6
Qwen2.5-Coder-14B 14B 80.5 91.0 88.5 87.7